Publications of the Astronomical Society of the Pacific Conference Series—Instructions for Authors Using L^{A}TEX Markup

Iman Author

The Name of My Institution, The Full Address of My Institution, Run in Address to Make Minimum Number of Lines

Ima Co-Author

The Name of My Institution, The Full Address of My Institution

Abstract. A concise abstract is recommended. Enter the text of the abstract in an “abstract” environment, i.e., within \begin{abstract} and \end{abstract} commands. Abstracts are required for all papers.

1. Introduction

In order to ensure that papers received for publication from different authors are consistent in format, style, and quality, authors are required to type their manuscripts exactly according to the following instructions. The editors will modify the electronic manuscripts as necessary to insure that they conform to these standards.

2. L^{A}TEX Markup Commands

Authors using L^{A}TEX are requested to use the paspconf style file, and declare it as a substyle to the standard L^{A}TEX article style. A copy of the style file, a sample paper, and further instructions are available via anonymous FTP.

You should use only those markup commands from L^{A}TEX plus the several extensions provided by this style file. Do not define any commands of your own for any reason (no \def or \newcommand statements).

2.1. Preamble

The first piece of markup in the manuscript must declare the overall style of the document.

     \documentstyle[11pt,paspconf]{article}

The \documentstyle command must appear first in any L^{A}TEX file, and this one specifies the main style to be the L^{A}TEX article style using eleven point fonts, with modifications and additions for the paspconf substyle.

The author must include a

     \begin{document}

command to identify the beginning of the main portion of the manuscript.

2.2. Title, Byline, Abstract, etc.

Title and author identification are by way of the standard L^{A}TEX commands \title and \author. An author’s principal affiliation is specified with a separate macro \affil. Each \author command should be followed by a corresponding \affil (address).

\title{lucid text}

     \author{name(s)}
     \affil{address}

The \affil command should be used to give the author’s full postal address. The address will be broken over several lines automatically; do not use L^{A}TEX’s \\ command to indicate the line breaks. Please use mixed case text for all these fields rather than supplying all capitals; the style file will convert to upper case as necessary.

The article’s abstract should be enclosed in an abstract environment.

     \begin{abstract}

abstract text

     \end{abstract}

Don’t include the word “Abstract” in your text; it is inserted automatically.

2.3. Sections

The L^{A}TEX article environment supports two levels of sectioning. (Actually, it supports more, but these are the relevant ones.)

     \section{heading}
     \subsection{heading}

Please use mixed case text for the section heads. Note that these commands delimit sections by marking the beginning of each section; there are not separate commands to identify the ends.

If one wishes to have an acknowledgments section, it should be set off simply with the command

     \acknowledgments

2.4. Equations

Displayed equations can be typeset in many ways using the standard displayed math environments of L^{A}TEX; these three are probably of greatest use.

     \begin{displaymath}
     \end{displaymath}
     \begin{equation}
     \end{equation}
     \begin{eqnarray}
     \end{eqnarray}

The displaymath environment will break out a single, unnumbered formula. The equation will appear the same if it is set in an equation environment, and it will be autonumbered by L^{A}TEX. In order to set several formulæ in which vertical alignment is required, use the eqnarray environment.

2.5. Tables

Tables should appear in table environments.

     \begin{table}
     \caption{text}
     \begin{tabular}{cols}
     \end{tabular}
     \end{table}

There should be only one table per environment. The table environment encloses not only the tabular material but also any title (caption) or footnote information associated with the table. Tabular information is typeset within L^{A}TEX’s tabular environment; the cols argument specifies the formatting for each column. Tables and figures will be identified with arabic numerals, e.g., “Table 2.”; the identifying text, including the number, is generated automatically by the \caption command.

There is a \tableline command for use in tabular environments.

     \tableline

This command produces a single horizontal rule. There should be a \tableline above and below between the column headings, and two at the end of the table. Authors should not use additional \tablelines themselves, and are discouraged from using vertical rules unless essential.

2.6. EPS Files

Authors who can prepare computer graphics in Encapsulated PostScript (EPS) format may use one of two additional markup commands to mark the point of inclusion, both of which should be used inside a L^{A}TEX figure environment. If the DVI translator dvips by Tom Rokicki is available on your computer, it is also possible to prepare the final copy with such figures in place. If you have Tom’s dvips program, you must include the epsf substyle in the \documentstyle definition:

     \documentstyle[11pt,paspconf,epsf]{article}

You may have to copy epsf.tex to epsf.sty or create a symbolic link, since dvips is distributed with epsf.tex which serves both plain TEX and L^{A}TEX. The same file works fine as a L^{A}TEX style file, it just needs the .sty extension.

The markup commands for including EPS files are:

     \plotone{file}
     \plottwo{file}{file}

The file argument is used to name the file(s) to be included. The \plotone command includes one figure that is scaled to the width of the current text column; \plottwo inserts two figures side by side, and the pair is scaled to fit the text width. If one uses these macros, the necessary vertical space is provided automatically.

     \begin{figure}

\plotone{mygraph.eps} \caption{My EPS graphic.}

     \end{figure}

or

     \begin{figure}

\plottwo{mygraph.eps}{another.eps} \caption{Two related graphics.}

     \end{figure}

Please note that the caption will be centered under the pair of graphics when \plottwo is used. It is not possible to caption the two plots individually with this package at this time. As with tables, figures will be identified with arabic numerals, e.g., “Figure 1.”.

If you must fuss with the positioning and scaling of the EPS plot on the printed page, you can try using this command:

\plotfiddle{file}{vsize}{rot}{hsf }{vsf }{htrans}{vtrans}

vsize vertical white space to allow for plot, any valid

L^{A}TEX dimension

rot  rotation angle, in degrees, counter-clockwise
hsf  horiz scale factor, percent
vsf  vert scale factor, percent
htrans horiz translation, in PS points 72/in
vtrans vert translation, in PS points 72/in

If you can produce EPS but you do not have dvips, you can still put the \plotone or \plottwo commands in the the appropriate places, but you will have to comment them out and put in a \vspace{dimen} command to open up the text. The dvips program is in the public domain and is available from labrea.stanford.edu.

A special note to authors: Color EPS files should be avoided if possible. And since it is sometimes necessary to edit EPS files to make them printable, authors should try to avoid EPS files with lines longer than 1024 characters.

2.7. Pasted in Illustrations

Illustrations must be inserted in the text at the appropriate places, with the relevant caption underneath each. The finished pages are reduced by 10% before printing. Thus, illustrations will appear somewhat smaller in print.

These illustrations should appear in figure environments.

     \begin{figure}
     \vspace{dimen}
     \caption{text}
     \end{figure}

There should be only one figure per environment. Space for the figure is created with the \vspace command; dimen should be a valid L^{A}TEX dimension, e.g.,

“2.5in”.

Stick glue is normally preferred for pasting in illustrations, but other adhesives will suffice. Do not use a glue which will wrinkle the paper. Illustrations must not be fixed to the pages by transparent adhesive (cellophane) tape since this will show in the printing. Paste in each illustration so that it is centered on the page. Also check (e.g., with a set-square) that the illustration is not crooked.

Size of Illustrations The maximum width of an illustration is normally 13.4cm (5.25in) so that it will fit within the width of the text area. Of course an illustration may be smaller if appropriate. A large illustration may be placed sideward (“landscape”) on the paper if necessary.

Halftone Illustrations (Photographs) Good glossy original prints are required, black and white only; color plates cannot be reproduced. Photographs cut from other publications will not reproduce well, and usually infringe copyright. Supply both a large photograph and a reduced version pasted in the text. The pasted-in version will be used for size and placement only; a copy machine can be used for this reduction. If a reduction is not supplied, leave appropriate space in text above the figure caption. The publisher will take care of the photographic reduction and mounting of the original glossy print in the space provided by your pasted-in version.

Line Illustrations If line drawings are submitted as the original, they should be drawn in black ink on good quality tracing paper or a glossy photographic print.

The lettering on line illustrations must be satisfactory — not too large, not too small. The preferred style of lettering is “all capitals”, with capital letters between 2mm and 3mm in height (roughly, 6pt to 8pt). Lettering should be done with a lettering stencil or press-on lettering; free-hand or typewritten lettering is not recommended. All lettering and fine detail on both line and photographic illustrations must be very clear and of good visual quality. It is often easier to prepare line illustrations at a larger size (e.g., twice the required size), then have them reduced photographically for pasting into the text.

Illustrations produced on laser printers are acceptable and are probably the preferred way to produce drawings by most authors.

3. References

3.1. In the Text

The reference system to be followed is the standard PASP or ApJ system: author name(s) followed by the year in parentheses, as in Abt (1990), or author and year both in parentheses (Abt 1990).

3.2. Reference List

There is a references environment that sets off the list of references and adjusts spacing parameters.

     \begin{references}

\reference bibliographic information

        .
        .
     \end{references}

The bibliographic information should be in the order directed by Abt (1990): author, year, journal, volume, and page. For instance, the reference for this editorial would be typed in as

     Abt, H. 1990, ApJ, 357, 1

Note that there is not a comma following the author name(s), there is not a trailing period at the end of the reference, and the entire line is set in the body typeface (no font changes).

Care should be taken that each literature citation in the manuscript has its counterpart in the reference list and vice versa. Care should also be given to checking the accuracy of the references—author(s), date, volume, and page number. The accuracy of the references is the sole responsibility of the author.

3.3. Abbreviations for Journals

There are macros for many of the oft-referenced journals so that authors may use the L^{A}TEX names rather than having to look up a particular journal’s specific abbreviation. Any stylistic requirements of the editors are taken care of by the macros, so authors need not be concerned about such editorial preferences.

\aj  Astronomical Journal
\araa Annual Review of Astronomy and Astrophysics
\apj  Astrophysical Journal
\apjl  , Letters to the Editor
\apjs  , Supplement Series
\ao  Applied Optics
\apss Astrophysics and Space Science
\aap  Astronomy and Astrophysics
\aaps  , Supplement Series
\azh  Astronomicheskii Zhurnal
\baas Bulletin of the AAS
\jrasc Journal of the RAS of Canada

\memras Memoirs of the RAS
\mnras Monthly Notices of the RAS
\pra  Physical Review A: General Physics
\prb  Physical Review B: Solid State
\prc  Physical Review C:
\prd  Physical Review D:
\prl  Physical Review Letters
\pasp  Publications of the ASP
\pasj  Publications of the ASJ
\qjras Quarterly Journal of the RAS
\skytel Sky and Telescope
\sovast Soviet Astronomy
\ssr  Space Science Reviews
\zap  Zeitschrift für Astrophysik

4. Examples

These instructions give an overview of the basic markup commands that need to be entered in a paper. Authors are encouraged to examine the sample papers that are included with the style file; these examples are named psample1.tex and psample2.tex. The file psample1.tex is a paper prepared with the PASP macros utilizing a minimal amount of markup. A more “complete” paper requiring most of the capabilities of the package is provided as psample2.tex; this file is annotated with comments that describe the purpose of the markup.

5. Final Check

Be sure you have read the README.<yourvol> file for any special instructions that may need to be followed for your proceedings volume.

6. Reprints

Reprints of papers for these Proceedings are not available. Articles may be copied from the published volume.

7. Submission of Manuscripts

See the README.<yourvol> file for instructions on submission of the manuscripts.